# Cherry picking in forests: A new characterization for the unrooted hybrid number of two phylogenetic trees

### v3

@@ 2. Preliminaries
Suppose that ${\mathcal{F}}$ is a forest on $X$. Then we say that ${\mathcal{N}}$ displays ${\mathcal{F}}$ if every tree in ${\mathcal{F}}$ is displayed by ${\mathcal{N}}$, and we can choose an image ${\mathcal{N}}[{\mathcal{T}}]$ for each tree ${\mathcal{T}}$ in ${\mathcal{F}}$ such that for any distinct trees ${\mathcal{T}},{\mathcal{T}}^{\prime}\in{\mathcal{F}}$ the images ${\mathcal{N}}[{\mathcal{T}}]$ and ${\mathcal{N}}[{\mathcal{T}}^{\prime}]$ do not share a vertex. If ${\mathcal{N}}$ displays ${\mathcal{F}}$, then we refer to the set ${\mathcal{N}}[{\mathcal{F}}]=\{{\mathcal{N}}[{\mathcal{T}}]\,:\,{\mathcal{T}}\in{\mathcal{F}}\}$ as an image of ${\mathcal{F}}$ in ${\mathcal{N}}$. For example, considering again the forest ${\mathcal{F}}^{\prime}$ and the phylogenetic network ${\mathcal{N}}$ pictured in Figure 1, the subgraph ${\mathcal{N}}[{\mathcal{F}}^{\prime}]$ of ${\mathcal{N}}$ is indicated in bold.
3. Cherry picking sequences
@@
Exploiting this observation, we shall now define a cherry picking sequence for two forests ${\mathcal{F}}$ and ${\mathcal{F}}^{\prime}$ on the same leaf set $X$ as follows: We call a sequence $\sigma=(x_{1},x_{2},\dots,x_{m})$, $m\geq|X|\geq 1$, of elements in $X$ a cherry picking sequence for ${\mathcal{F}}$ and ${\mathcal{F}}^{\prime}$ (of length $m$) if there are forests ${\mathcal{F}}[i]$ and ${\mathcal{F}}^{\prime}[i]$, $1\leq i\leq m$, such that ${\mathcal{F}}[1]={\mathcal{F}}$, ${\mathcal{F}}^{\prime}[1]={\mathcal{F}}^{\prime}$, ${\mathcal{F}}[m]={\mathcal{F}}^{\prime}[m]=\{x_{m}\}$ and, for each $11\leq i\leq m-$ 1, precisely one of the following cases holds. The different cases are illustrated in Figure 2 and an example of a cherry picking sequence is presented in Figure 3.
@@
Figure 2. The different cases as described in the definition of a cherry picking sequence. Ovals indicate subtrees. The roles of ${\mathcal{F}}$ and ${\mathcal{F}}^{\prime}$ in (C2) and (C3) could also be reversed.
Refer to caption222255556666222244446666222266662222555566662222444455556666111133332222444455556666555566662222444422224444333311115555666655556666222222222222333311113333111155556666555566662222444422224444333333335555666622224444apply (C2)(a)(i) to $x_{1}=1$${\mathcal{F}}[3]$${\mathcal{F}}^{\prime}[3]$apply (C3) to $x_{2}=1$${\mathcal{F}}^{\prime}[6]$apply (C1) to $x_{6}=6$${\mathcal{F}}={\mathcal{F}}[1]$${\mathcal{F}}[5]$${\mathcal{F}}^{\prime}[5]$${\mathcal{F}}[7]$${\mathcal{F}}[2]$${\mathcal{F}}^{\prime}[2]$${\mathcal{F}}^{\prime}[7]$apply (C3) to $x_{3}=3$${\mathcal{F}}[4]$${\mathcal{F}}^{\prime}={\mathcal{F}}^{\prime}[1]$${\mathcal{F}}^{\prime}[4]$apply (C1) to $x_{4}=4$apply (C1) to $x_{5}=5$${\mathcal{F}}[6]$
Figure 3. For the two depicted forests ${\mathcal{F}}$ and ${\mathcal{F}}^{\prime}$ the sequence $(1,1,3,4,5,6,2)$ is a cherry picking sequence for ${\mathcal{F}}$ and ${\mathcal{F}}^{\prime}$. The reductions applied within the sequence are indicated to the left of the arrows.
@@
Let $\sigma=(x_{1},\dots,x_{m})$ be a cherry picking sequence for two forests ${\mathcal{F}}$ and ${\mathcal{F}}^{\prime}$ with $m\geq 2$. For each $11\leq i\leq m-$ 1, we set $c(x_{i})=0$ if one of the reductions in (C1) and (C3) is applied to $x_{i}$ and, otherwise, $c(x_{i})=1$. In addition, we define the weight of $\sigma$ by $w(\sigma)=\sum_{i=1}^{m-1}c(x_{i})$ if $m\geq 2$ and 00 if $m=1$. To illustrate these definitions, let $X=\{1,\ldots,6\}$ and consider the two forests ${\mathcal{F}}$ and ${\mathcal{F}}^{\prime}$ on $X$ in Figure 1. Then $(1,1,3,4,5,6,2)$ gives a cherry picking sequence for ${\mathcal{F}}$ and ${\mathcal{F}}^{\prime}$ with weight 1, as can be verified by looking at the reductions depicted in Figure 3. We now show that a cherry picking sequence always exists for any pair of forests with the same leaf set.
Proposition 3.2.
@@
For case (C2), since one of the cases (a), (b) or (c) must hold, we can apply the reduction ${\mathcal{F}}_{1}={\mathcal{F}}[i]-e_{x}$ and ${\mathcal{F}}_{2}={\mathcal{F}}^{\prime}[i]$ (which occurs in cases (a)(i), (b)(i) and (c)). This implies that $x$ is a component in ${\mathcal{F}}_{1}$. So, we can then apply a (C3)-reduction to ${\mathcal{F}}_{1}$ and ${\mathcal{F}}_{2}$ with $x_{i}=x$ which results in two forests ${\mathcal{F}}_{1}^{\prime\prime}$ and ${\mathcal{F}}_{2}^{\prime\prime}$ on $X\setminus\{x\}$. By induction, there is a cherry picking sequence $(x_{1},\dots,x_{p})$ for ${\mathcal{F}}_{1}^{\prime\prime}$ and ${\mathcal{F}}_{2}^{\prime\prime}$ with $p\geq n-1$. It therefore follows that $(x,x,x_{1},\dots,x_{p})$ is a cherry picking sequence for ${\mathcal{F}}$ and ${\mathcal{F}}^{\prime}$ with length $p+2>n$. ∎
@@
4. Bounding the hybrid number from above using cherry picking sequences
@@
Before proceeding, denoting by $d_{TBR}({\mathcal{T}},{\mathcal{T}}^{\prime})$ the TBR distance between two phylogenetic trees ${\mathcal{T}}$ and ${\mathcal{T}}^{\prime}$ on $X$ [1, 25], we note that as an immediate corollary of Theorem 4.1 and [13, Theorem 3] we obtain:
@@
Suppose that ${\mathcal{F}}$ and ${\mathcal{F}}^{\prime}$ are forests on $X$, $|X|\geq 2$, that either (i) both ${\mathcal{F}}$ and ${\mathcal{F}}^{\prime}$ contain a cherry $(x,y)$, $x,y\in X$, or (ii) ${\mathcal{F}}$ contains a component $x$, $x\in X$. If ${\mathcal{N}}^{\prime}$ is a phylogenetic network on $X\backslash\{x\}$ that displays ${\mathcal{F}}-x$ and ${\mathcal{F}}^{\prime}-x$, then there is a phylogenetic network ${\mathcal{N}}$ on $X$ with $r({\mathcal{N}})=r({\mathcal{N}}^{\prime})$ that displays ${\mathcal{F}}$ and ${\mathcal{F}}^{\prime}$.
Proof.
@@
5. Four technical lemmas
In this section, we shall prove four technical lemmas (Lemmas 5.2, 5.4, 5.5 and 5.6) that are concerned with how we can reduce the reticulation number of a phylogenetic network displaying two forests whilst retaining the displaying property. We shall use these lemmas to prove Theorem 6.1 in the next section.
@@ 5. Four technical lemmas
Now, let ${\mathcal{N}}$ be a pseudo-network on $X$. A maximal 2-connected component of ${\mathcal{N}}$ that is not a single edge is called a blob of ${\mathcal{N}}$. Let ${\mathcal{B}}$ be a blob of ${\mathcal{N}}$, and let $e=\{u,v\}$ be an edge of ${\mathcal{N}}$. We say that ${\mathcal{B}}$ is incident with $e$ (or that $e$ is incident with ${\mathcal{B}}$) if precisely one of $u$ and $v$ is contained in $V({\mathcal{B}})$. Furthermore, if $e$ is incident with ${\mathcal{B}}$ and contains a leaf $x\in X$, then we say that ${\mathcal{B}}$ is incident with $x$. We denote the subset of leaves of ${\mathcal{N}}$ that are incident with ${\mathcal{B}}$ by $L({\mathcal{B}})$. Now let $E$ be the subset of edges of ${\mathcal{N}}$ that are incident with ${\mathcal{B}}$. We refer to ${\mathcal{B}}$ as a pendant blob of ${\mathcal{N}}$ if at most one edge in $E$ is a non-trivial cut-edge of ${\mathcal{N}}$ and all other edges in $E$ are incident with a leaf.
Let ${\mathcal{N}}$ be a pseudo-network that displays two forests ${\mathcal{F}}$ and ${\mathcal{F}}^{\prime}$. In what follows, we shall be interested in understanding how $r({\mathcal{N}})$ changes if we transform ${\mathcal{N}}$ into a phylogenetic network that displays ${\mathcal{F}}$ and ${\mathcal{F}}^{\prime}$ by removing certain leaves or edges. To this end, suppose that ${\mathcal{N}}$ is a pseudo-network on $X$, $|X|\geq 2$, with precisely one multi-edge. We shall call a sequence ${\mathcal{N}}_{1},\ldots,{\mathcal{N}}_{k}$, $k\geq 2$, of distinct pseudo-networks on $X$ a simplification sequence for ${\mathcal{N}}$ if ${\mathcal{N}}={\mathcal{N}}_{1}$, each ${\mathcal{N}}_{i+1}$ is obtained from ${\mathcal{N}}_{i}$ by suppressing a single multi-edge in ${\mathcal{N}}_{i}$ and suppressing the resulting vertices of degree two for all $11\leq i\leq k-$ 1, and ${\mathcal{N}}_{k}$ is a phylogenetic network on $X$.
To prove the next lemma, we will use the following observation whose proof is straight-forward.
Observation 5.1.
Suppose that ${\mathcal{N}}=(V,E)$ is a pseudo-network with a unique multi-edge between two vertices $u$ and $v$ in $V$, and that $\{v,p\}$ and $\{u,q\}$ are edges of ${\mathcal{N}}$ with $p,q\in V\setminus\{u,v\}$. Then the graph obtained by suppressing the multi-edge and resulting degree 2 vertices is a pseudo-network if and only if $p\not=q$.
@@
Suppose that ${\mathcal{F}}$ and ${\mathcal{F}}^{\prime}$ are two forests on $X$ with $|X|\geq 2$ and that ${\mathcal{N}}$ is a pseudo-network on $X$ with precisely one multi-edge such that ${\mathcal{N}}$ displays ${\mathcal{F}}$ and ${\mathcal{F}}^{\prime}$. In addition, suppose that the blob in ${\mathcal{N}}$ that contains the multi-edge is incident with at least two cut-edges of ${\mathcal{N}}$. Then there exists a simplification sequence ${\mathcal{N}}_{1},\ldots,{\mathcal{N}}_{k}$ for ${\mathcal{N}}$, some $k\geq 2$. Moreover, ${\mathcal{N}}_{k}$ displays ${\mathcal{F}}$ and ${\mathcal{F}}^{\prime}$ and $r({\mathcal{N}})>r({\mathcal{N}}_{k})$.
Proof.
Since ${\mathcal{N}}$ has precisely one multi-edge and the blob that contains this multi-edge is incident with at least two cut-edges of ${\mathcal{N}}$, by applying Observation 5.1 it is straight-forward to check that the multi-graph ${\mathcal{N}}^{\prime}$ obtained by suppressing the unique multi-edge and any resulting degree-two vertex has a blob ${\mathcal{B}}^{\prime}$ that is incident with at least two cut-edges, and is either a phylogenetic network on $X$, or it is a pseudo-network on $X$ with a unique multi-edge that is contained in ${\mathcal{B}}^{\prime}$. This implies immediately that there is a unique simplification sequence ${\mathcal{N}}_{1},\ldots,{\mathcal{N}}_{k}$ for ${\mathcal{N}}$, some $k\geq 2$. We now show that ${\mathcal{N}}_{i}$ displays the forests ${\mathcal{F}}$ and ${\mathcal{F}}^{\prime}$, for all $1\leq i\leq k$. For some $i\in\{1,\ldots,k-1\}$, suppose that ${\mathcal{N}}_{i}$ displays the forests ${\mathcal{F}}$ and ${\mathcal{F}}^{\prime}$, that $e=\{u,v\}$ and $e^{\prime}=\{u,v\}$ are edges of ${\mathcal{N}}_{i}$ and that, $e$ is suppressed to obtain ${\mathcal{N}}_{i+1}$. Choose images for ${\mathcal{F}}$ and ${\mathcal{F}}^{\prime}$ in ${\mathcal{N}}$, respectively. It is straight-forward to check that if $e$ is not contained in the image of ${\mathcal{F}}$ or ${\mathcal{F}}^{\prime}$, then the pseudo-network ${\mathcal{N}}_{i+1}$ displays ${\mathcal{F}}$ and ${\mathcal{F}}^{\prime}$. Moreover, if one of $e$ and $e^{\prime}$, say $e$, is contained in the image of ${\mathcal{F}}$ and $e^{\prime}$ is contained in the image of ${\mathcal{F}}^{\prime}$, then clearly we can alter the image of ${\mathcal{F}}$ so that only $e^{\prime}$ is contained in the image of ${\mathcal{F}}$ and ${\mathcal{F}}^{\prime}$. Since ${\mathcal{N}}$ displays ${\mathcal{F}}$ and ${\mathcal{F}}^{\prime}$, it follows that ${\mathcal{N}}_{i+1}$ also does for all $11\leq i\leq k-$ 1.
Finally, we note that $r({\mathcal{N}})>r({\mathcal{N}}_{k})$ since clearly $r({\mathcal{N}}_{i})=r({\mathcal{N}}_{i+1})+1$, for all $11\leq i\leq k-$ 1. ∎
@@
Figure 4. An explicit example of Lemma 5.4 applied to $x=3$ of a phylogenetic network ${\mathcal{N}}$ on $\{1,2,\ldots,5\}$ with a pendant blob ${\mathcal{B}}$ that displays two forests ${\mathcal{F}}$ and ${\mathcal{F}}^{\prime}$ on $\{1,2\ldots,5\}$. Then ${\mathcal{N}}-3$ is not a phylogenetic network, and there exists a phylogenetic network ${\mathcal{N}}^{\prime}$ on $\{1,2,4,5\}$ with $r({\mathcal{N}})>r({\mathcal{N}}^{\prime})$ that displays ${\mathcal{F}}-3$ and ${\mathcal{F}}^{\prime}-3$.
Using Observation 5.3, we obtain:
@@
Proof.
Suppose first that $X=\{x,y\}$. Then we can take ${\mathcal{N}}^{\prime}$ to be the vertex $y$ to complete the proof of the lemma. Now, suppose that $|X|\geq 3$. Note that ${\mathcal{N}}-x$ clearly displays ${\mathcal{F}}-x$ and ${\mathcal{F}}^{\prime}-x$. Let $v$ and $w$ be the two vertices adjacent to the non-leaf vertex of $e_{x}$. Since ${\mathcal{N}}$ is a phylogenetic network, we have $v\neq w$. Moreover, by Observation 5.3, $\{v,w\}$ is an edge of ${\mathcal{N}}$ and ${\mathcal{N}}-x$ contains the unique multi-edge between $v$ and $w$. This multi-edge is contained in a blob ${\mathcal{B}}^{\prime}$ of ${\mathcal{N}}-x$. Since ${\mathcal{B}}$ is a pendant blob in ${\mathcal{N}}$, it follows that ${\mathcal{B}}^{\prime}$ is a pendant blob in ${\mathcal{N}}-x$. Furthermore, since $|L({\mathcal{B}})|\geq 2$ and $|X|\geq 3$, either $|L({\mathcal{B}}^{\prime})|\geq 2$, or $|L({\mathcal{B}}^{\prime})|=1$ and ${\mathcal{B}}^{\prime}$ is incident with at least one non-trivial cut-edge. The proof of the lemma now follows immediately by applying Lemma 5.2 to ${\mathcal{N}}-x$ and ${\mathcal{B}}^{\prime}$. ∎
@@
If ${\mathcal{N}}^{\prime}$ is a phylogenetic network, this completes the proof of the lemma. Otherwise, since by assumption ${\mathcal{B}}$ must be incident with at least two cut-edges, it follows that we can apply Lemma 5.2 to complete the proof. ∎
@@
  • Suppose $|L({\mathcal{B}})|=1$, so that $L({\mathcal{B}})=\{x\}$, some $x\in X$. If ${\mathcal{N}}={\mathcal{B}}$, then define ${\mathcal{N}}-{\mathcal{B}}=\{x\}$. Otherwise, we let ${\mathcal{N}}-{\mathcal{B}}$ be the phylogenetic network obtained by first removing all vertices in ${\mathcal{B}}$ and edges that contain them from ${\mathcal{N}}$ and then adding the edge $\{x,u\}$ (i.e. the graph obtained by replacing ${\mathcal{B}}$ with leaf $x$).
@@
6. Bounding the weight of a cherry picking sequence from above by the hybrid number

### v5

@@ 2 Preliminaries
Suppose that ${\mathcal{F}}$ is a forest on $X$. Then we say that ${\mathcal{N}}$ displays ${\mathcal{F}}$ if every tree in ${\mathcal{F}}$ is displayed by ${\mathcal{N}}$, and we can choose an image ${\mathcal{N}}[{\mathcal{T}}]$ for each tree ${\mathcal{T}}$ in ${\mathcal{F}}$ such that for any distinct trees ${\mathcal{T}},{\mathcal{T}}^{\prime}\in{\mathcal{F}}$ the images ${\mathcal{N}}[{\mathcal{T}}]$ and ${\mathcal{N}}[{\mathcal{T}}^{\prime}]$ do not share a vertex. If ${\mathcal{N}}$ displays ${\mathcal{F}}$, then we refer to the set ${\mathcal{N}}[{\mathcal{F}}]=\{{\mathcal{N}}[{\mathcal{T}}]\,:\,{\mathcal{T}}\in{\mathcal{F}}\}$ as an image of ${\mathcal{F}}$ in ${\mathcal{N}}$. For example, considering again the forest ${\mathcal{F}}^{\prime}$ and the phylogenetic network ${\mathcal{N}}$ pictured in Fig. 1, the subgraph ${\mathcal{N}}[{\mathcal{F}}^{\prime}]$ of ${\mathcal{N}}$ is indicated in bold.
3 Cherry picking sequences
[…]
Figure 2: The different cases as described in the definition of a cherry picking sequence. Ovals indicate subtrees. The roles of ${\mathcal{F}}$ and ${\mathcal{F}}^{\prime}$ in (C2) and (C3) could also be reversed.
@@ 3 Cherry picking sequences
Exploiting this observation, we shall now define a cherry picking sequence for two forests ${\mathcal{F}}$ and ${\mathcal{F}}^{\prime}$ on the same leaf set $X$ as follows: We call a sequence $\sigma=(x_{1},x_{2},\dots,x_{m})$, $m\geq|X|\geq 1$, of elements in $X$ a cherry picking sequence for ${\mathcal{F}}$ and ${\mathcal{F}}^{\prime}$ (of length $m$) if there are forests ${\mathcal{F}}[i]$ and ${\mathcal{F}}^{\prime}[i]$, $1\leq i\leq m$, such that ${\mathcal{F}}[1]={\mathcal{F}}$, ${\mathcal{F}}^{\prime}[1]={\mathcal{F}}^{\prime}$, ${\mathcal{F}}[m]={\mathcal{F}}^{\prime}[m]=\{x_{m}\}$ and, for each $11\leq i\leq m-$ 1, precisely one of the following cases holds. The different cases are illustrated in Fig. 2 and an example of a cherry picking sequence is presented in Fig. 3.
@@
Figure 3: For the two depicted forests ${\mathcal{F}}$ and ${\mathcal{F}}^{\prime}$ the sequence $(8,1,9,10,8,8,2,6,7,1,1,3,4,5,6,2)$ is a cherry picking sequence for ${\mathcal{F}}$ and ${\mathcal{F}}^{\prime}$. The reductions applied within the sequence are indicated next to the arrows.
@@
Let $\sigma=(x_{1},\dots,x_{m})$ be a cherry picking sequence for two forests ${\mathcal{F}}$ and ${\mathcal{F}}^{\prime}$ with $m\geq 2$. For each $11\leq i\leq m-$ 1, we set $c(x_{i})=0$ if one of the reductions in (C1) and (C3) is applied to $x_{i}$ and, otherwise, $c(x_{i})=1$. In addition, we define the weight of $\sigma$ by $w(\sigma)=\sum_{i=1}^{m-1}c(x_{i})$ if $m\geq 2$ and 00 if $m=1$. To illustrate these definitions, let $X=\{1,\ldots,6\}$ and consider the two forests ${\mathcal{F}}$ and ${\mathcal{F}}^{\prime}$ on $X$ in Fig. 1. Then $(8,1,9,10,8,8,2,6,7,1,1,3,4,5,6,2)$ gives a cherry picking sequence for ${\mathcal{F}}$ and ${\mathcal{F}}^{\prime}$ with weight 6, as can be verified by looking at the reductions depicted in Fig. 3.
We now show that a cherry picking sequence always exists for any pair of forests with the same leaf set.
Proposition 3.2
@@
For case (C2), since one of the cases (a), (b) or (c) must hold, we can apply the reduction ${\mathcal{F}}_{1}={\mathcal{F}}[i]-e_{x}$ and ${\mathcal{F}}_{2}={\mathcal{F}}^{\prime}[i]$ (which occurs in cases (a)(i), (b)(i) and (c)). This implies that $x$ is a component in ${\mathcal{F}}_{1}$. So, we can then apply a (C3)-reduction to ${\mathcal{F}}_{1}$ and ${\mathcal{F}}_{2}$ with $x_{i}=x$ which results in two forests ${\mathcal{F}}_{1}^{\prime\prime}$ and ${\mathcal{F}}_{2}^{\prime\prime}$ on $X\setminus\{x\}$. By induction, there is a cherry picking sequence $(x_{1},\dots,x_{p})$ for ${\mathcal{F}}_{1}^{\prime\prime}$ and ${\mathcal{F}}_{2}^{\prime\prime}$ with $p\geq n-1$. It now follows that $(x,x,x_{1},\dots,x_{p})$ is a cherry picking sequence for ${\mathcal{F}}$ and ${\mathcal{F}}^{\prime}$ with length $p+2>n$.
@@
4 Bounding the hybrid number from above using cherry picking sequences
@@
Before proceeding, denoting by $d_{TBR}({\mathcal{T}},{\mathcal{T}}^{\prime})$ the TBR distance between two phylogenetic trees ${\mathcal{T}}$ and ${\mathcal{T}}^{\prime}$ on $X$ [1, 26], we note that as an immediate corollary of Theorem 4.1 and [14, Theorem 3] we obtain:
@@
Suppose that ${\mathcal{F}}$ and ${\mathcal{F}}^{\prime}$ are forests on $X$, $|X|\geq 2$, such that either (i) both ${\mathcal{F}}$ and ${\mathcal{F}}^{\prime}$ contain a cherry $(x,y)$, $x,y\in X$, or (ii) ${\mathcal{F}}$ contains a component $x$, $x\in X$. If ${\mathcal{N}}^{\prime}$ is a phylogenetic network on $X\backslash\{x\}$ that displays ${\mathcal{F}}-x$ and ${\mathcal{F}}^{\prime}-x$, then there is a phylogenetic network ${\mathcal{N}}$ on $X$ with $r({\mathcal{N}})=r({\mathcal{N}}^{\prime})$ that displays ${\mathcal{F}}$ and ${\mathcal{F}}^{\prime}$.
Proof 4.4.
@@
We end this section by noting that the proof of Theorem 4.7 in combination with the constructive proofs of Lemmas 4.3 and 4.5 lend themselves to a polynomial-time algorithm that reconstructs a phylogenetic network ${\mathcal{N}}$ from a given cherry picking sequence $\sigma$ such that $w(\sigma)\geq r({\mathcal{N}})$.
5 Four technical lemmas
In this section, we shall prove four technical lemmas (Lemmas 5.1, 5.4, 5.6 and 5.8) that are concerned with how we can reduce the reticulation number of a phylogenetic network displaying two forests whilst retaining the displaying property. We shall use these lemmas to prove Theorem 6.1 in the next section.
@@ 5 Four technical lemmas
Now, let ${\mathcal{N}}$ be a pseudo-network on $X$. A maximal 2-connected [^1] A graph $G$ is 2-connected if the graph obtained by removing any vertex from $G$ is connected. subgraph of ${\mathcal{N}}$ that is not an edge is called a blob of ${\mathcal{N}}$. In particular, note that a multi-edge in ${\mathcal{N}}$ could be a blob. Let ${\mathcal{B}}$ be a blob of ${\mathcal{N}}$, and let $e=\{u,v\}$ be an edge of ${\mathcal{N}}$. We say that ${\mathcal{B}}$ is incident with $e$ (or that $e$ is incident with ${\mathcal{B}}$) if precisely one of $u$ and $v$ is contained in $V({\mathcal{B}})$. Furthermore, if $e$ is incident with ${\mathcal{B}}$ and contains a leaf $x\in X$, then we say that ${\mathcal{B}}$ is incident with $x$. We denote the subset of leaves of ${\mathcal{N}}$ that are incident with ${\mathcal{B}}$ by $L({\mathcal{B}})$. Now let $E$ be the subset of edges of ${\mathcal{N}}$ that are incident with ${\mathcal{B}}$. We refer to ${\mathcal{B}}$ as a pendant blob of ${\mathcal{N}}$ if at most one edge in $E$ is a non-trivial cut-edge of ${\mathcal{N}}$ and all other edges in $E$ are incident with a leaf.
In what follows, for a pseudo-network ${\mathcal{N}}$ on $X$ that displays two forests ${\mathcal{F}}$ and ${\mathcal{F}}^{\prime}$, we shall be interested in understanding how $r({\mathcal{N}})$ changes if we transform ${\mathcal{N}}$ into a phylogenetic network that displays ${\mathcal{F}}$ and ${\mathcal{F}}^{\prime}$ by removing certain leaves or edges. To this end, suppose that $|X|\geq 2$, and ${\mathcal{N}}$ contains precisely one multi-edge. We shall call a sequence ${\mathcal{N}}_{1},\ldots,{\mathcal{N}}_{k}$, $k\geq 2$, of distinct pseudo-networks on $X$ a simplification sequence for ${\mathcal{N}}$ if ${\mathcal{N}}={\mathcal{N}}_{1}$, each ${\mathcal{N}}_{i+1}$ is obtained from ${\mathcal{N}}_{i}$ by suppressing a single multi-edge in ${\mathcal{N}}_{i}$ and suppressing the resulting vertices of degree two for all $11\leq i\leq k-$ 1, and ${\mathcal{N}}_{k}$ is a phylogenetic network on $X$.
@@
Suppose that ${\mathcal{F}}$ and ${\mathcal{F}}^{\prime}$ are two forests on $X$ with $|X|\geq 2$ and that ${\mathcal{N}}$ is a pseudo-network on $X$ with precisely one multi-edge such that ${\mathcal{N}}$ displays ${\mathcal{F}}$ and ${\mathcal{F}}^{\prime}$. In addition, suppose that the blob in ${\mathcal{N}}$ that contains the multi-edge is incident with at least two cut-edges of ${\mathcal{N}}$. Then there exists a simplification sequence ${\mathcal{N}}_{1},\ldots,{\mathcal{N}}_{k}$ for ${\mathcal{N}}$, for some $k\geq 2$. Moreover, ${\mathcal{N}}_{k}$ displays ${\mathcal{F}}$ and ${\mathcal{F}}^{\prime}$ and $r({\mathcal{N}})>r({\mathcal{N}}_{k})$.
Proof 5.2.
Let ${\mathcal{B}}$ be the blob in ${\mathcal{N}}$ that contains the unique multi-edge in ${\mathcal{N}}$, and let $u$ and $v$ be the two vertices contained in the multi-edge. In addition let $p$ and $q$ be the vertices in ${\mathcal{N}}$ such that $\{u,p\}$ and $\{v,q\}$ are in $E({\mathcal{N}})$. Then since ${\mathcal{B}}$ is incident with at least two cut-edges it follows that $p\neq q$. Thus, the pseudo-network obtained by suppressing the multi-edge and any resulting degree-two vertex is either a phylogenetic network on $X$, or contains a blob that is incident with at least two cut-edges and contains a unique multi-edge. This immediately implies that there is a unique simplification sequence ${\mathcal{N}}_{1},\ldots,{\mathcal{N}}_{k}$ for ${\mathcal{N}}$, for some $k\geq 2$.
We now show that ${\mathcal{N}}_{i}$ displays the forests ${\mathcal{F}}$ and ${\mathcal{F}}^{\prime}$, for all $1\leq i\leq k$. For some $i\in\{1,\ldots,k-1\}$, suppose that ${\mathcal{N}}_{i}$ displays the forests ${\mathcal{F}}$ and ${\mathcal{F}}^{\prime}$, that $e=\{u,v\}$ and $e^{\prime}=\{u,v\}$ are edges of ${\mathcal{N}}_{i}$ and that, $e$ is suppressed to obtain ${\mathcal{N}}_{i+1}$. Choose images for ${\mathcal{F}}$ and ${\mathcal{F}}^{\prime}$ in ${\mathcal{N}}$, respectively. It is straight-forward to check that if $e$ is not contained in the image of ${\mathcal{F}}$ or ${\mathcal{F}}^{\prime}$, then the pseudo-network ${\mathcal{N}}_{i+1}$ displays ${\mathcal{F}}$ and ${\mathcal{F}}^{\prime}$. Moreover, if one of $e$ and $e^{\prime}$, say $e$, is contained in the image of ${\mathcal{F}}$ and $e^{\prime}$ is contained in the image of ${\mathcal{F}}^{\prime}$, then clearly we can alter the image of ${\mathcal{F}}$ so that only $e^{\prime}$ is contained in the image of ${\mathcal{F}}$ and ${\mathcal{F}}^{\prime}$. Since ${\mathcal{N}}$ displays ${\mathcal{F}}$ and ${\mathcal{F}}^{\prime}$, it follows that ${\mathcal{N}}_{i+1}$ also does for all $11\leq i\leq k-$ 1.
Finally, we note that $r({\mathcal{N}})>r({\mathcal{N}}_{k})$ since clearly $r({\mathcal{N}}_{i})=r({\mathcal{N}}_{i+1})+1$, for all $11\leq i\leq k-$ 1.
@@
Figure 4: An explicit example of Lemma 5.4 applied to $x=3$ of a phylogenetic network ${\mathcal{N}}$ on $\{1,2,\ldots,5\}$ with a pendant blob ${\mathcal{B}}$ that displays two forests ${\mathcal{F}}$ and ${\mathcal{F}}^{\prime}$ on $\{1,2\ldots,5\}$. Then ${\mathcal{N}}-3$ is not a phylogenetic network, and there exists a phylogenetic network ${\mathcal{N}}^{\prime}$ on $\{1,2,4,5\}$ with $r({\mathcal{N}})>r({\mathcal{N}}^{\prime})$ that displays ${\mathcal{F}}-3$ and ${\mathcal{F}}^{\prime}-3$.
Using Observation 5.3, we obtain the following lemma, which is illustrated in Fig. 4.
@@
Proof 5.5.
Suppose first that $X=\{x,y\}$. Then we can take ${\mathcal{N}}^{\prime}$ to be the vertex $y$ to complete the proof of the lemma.
Now, suppose that $|X|\geq 3$. Note that ${\mathcal{N}}-x$ clearly displays ${\mathcal{F}}-x$ and ${\mathcal{F}}^{\prime}-x$. Let $v$ and $w$ be the two vertices adjacent to the non-leaf vertex of $e_{x}$. Since ${\mathcal{N}}$ is a phylogenetic network, we have $v\neq w$. Moreover, by Observation 5.3, $\{v,w\}$ is an edge of ${\mathcal{N}}$ and ${\mathcal{N}}-x$ contains the unique multi-edge between $v$ and $w$. This multi-edge is contained in a blob ${\mathcal{B}}^{\prime}$ of ${\mathcal{N}}-x$. Since ${\mathcal{B}}$ is a pendant blob in ${\mathcal{N}}$, it follows that ${\mathcal{B}}^{\prime}$ is a pendant blob in ${\mathcal{N}}-x$. Furthermore, since $|L({\mathcal{B}})|\geq 2$ and $|X|\geq 3$, either $|L({\mathcal{B}}^{\prime})|\geq 2$, or $|L({\mathcal{B}}^{\prime})|=1$ and ${\mathcal{B}}^{\prime}$ is incident with at least one non-trivial cut-edge. The proof of the lemma now follows immediately by applying Lemma 5.1 to ${\mathcal{N}}-x$ and ${\mathcal{B}}^{\prime}$.
@@
If ${\mathcal{N}}^{\prime}$ is a phylogenetic network, this completes the proof of the lemma. Otherwise, since by assumption ${\mathcal{B}}$ must be incident with at least two cut-edges, it follows that we can apply Lemma 5.1 to complete the proof.
@@
  • Suppose $|L({\mathcal{B}})|=1$, so that $L({\mathcal{B}})=\{x\}$, for some $x\in X$. If ${\mathcal{N}}={\mathcal{B}}$, then define ${\mathcal{N}}-{\mathcal{B}}=\{x\}$. Otherwise, we let ${\mathcal{N}}-{\mathcal{B}}$ be the phylogenetic network obtained by first removing all vertices in ${\mathcal{B}}$ and edges that contain them from ${\mathcal{N}}$ and then adding the edge $\{x,u\}$ (i.e. the graph obtained by replacing ${\mathcal{B}}$ with leaf $x$).
@@
6 Bounding the weight of a cherry picking sequence from above by the hybrid number